OpenStudy (hayleymeyer):

The diameter of Earth is about 7,926 miles and the diameter of Mars is about 4,194 miles. How much more surface area is there on Earth than on Mars to the nearest hundred million square miles? please heeellp! xoxo

OpenStudy (anonymous):

\[S_E-S_M=4*\pi*R_E^2-4*\pi*R_M^2=4*\pi*(R_E^2-R_M^2)=4*3.14(3963^2-2094^2)=11307960*12.56=142027977\]